I need to understand how to solve this equation and to find the value of x.
1/2(2x-1)+1/4(x-2)=4
Thanks

Found 2 solutions by fabioGR, elima:
Answer by fabioGR(2)

4=4*4
2(2x-1)+(x-2)=16
4x-2+x-2=16
5x-4=16
5x=20
x=20/5
x=4

Answer by elima(1433)

1/2(2x-1)+1/4(x-2)=4
You begin by multiplying 1/2(2x-1)= 2x-1/2
then 1/4(x-2) = x-2/4
Now you have:
2x-1/2+x-2/4=4
To add the left side together you need common denominators,which would be 4.
So; 2(2x-1)/4+x-2/4=4
Now you can add the left side together:
4x-2/4+x-2/4 = 5x-4/4
5x-4/4=4
Now multiply 4 to both sides so you can get rid of the 4 denominator.
5x-4 = 16
add 4 to both sides so you can get 5x by itself.
5x=20
Divide both sides by 5;
x=4